an open organ pipe has afrequency f. if half part of the organ pipe is dipped in water then what will be the change in frequency?

Grade:

2 Answers

Vbhootna Vb
11 years ago

freq. remain same

Harish R
11 years ago

When you place the pipe in water the organ pipe acts as a closed one. 

Initially frequency of the organ pipe is f = V/2L

When you place in water with only half the length,

the frequency becomes f = V/4L/2 = V/2L 

Hence frequency doesnt change

If you like this answer please approve it by clicking on the "Yes" given below!!!!Smile